# Concentration and fluctuations of sine–Gordon measure around topological multi-soliton manifold

### v1

@@ 3.2. Multi-topological solitons
$=-\frac{1}{2}\sum_{e}\Big(\cos((1-e_{1})m_{i_{1}}+\cdots+(1-e_{n})m_{i_{n}})+\cos((1+e_{1})m_{i_{1}}+\cdots+(1+e_{n})m_{i_{n}})\Big)\prod_{j=1}^{n}e_{j}$
@@
Before concluding this subsection, we present the following quantified energy gap estimate for the middle-range case.
Lemma 3.13.
Let $Q\in\mathbb{Z}$ with $Q\geq 2$, and let $d>0$ be a large constant. Then there exists $c>0$ such that if $\phi\in\mathcal{C}_{Q}$ with $\phi(-\infty)=0$ and $\phi(\infty)=2\pi Q$ satisfies
$\textup{dist}(\phi,\mathcal{M}_{Q}^{d\leq\cdot<(1+\eta)d}):=\inf_{\begin{subarray}{c}\xi_{1},\dots,\xi_{Q}\in\mathbb{R}\\ d\leq\min_{i\neq j}|\xi_{i}-\xi_{j}|<(1+\eta)d\end{subarray}}\|\phi-m_{\xi_{1},\dots,\xi_{Q}}\|_{L^{2}(\mathbb{R})}<\delta$ (3.21)
where $\frac{1}{2}e^{-d}\geq c\cdot\delta$, then
$E(\phi)\geq|Q|E_{\textup{kink}}+c\cdot e^{-(1+\eta)d}.$
Proof.
The proof is almost the same as that of Lemma 3.10. We only present the different parts. By following (3.11) and (3.15), we can find $\xi_{1}^{*},\dots,\xi_{Q}^{*}$ such that
$d\leq\min_{i\neq j}|\xi_{i}^{*}-\xi_{j}^{*}|<(1+\eta)d\quad\textup{and}\quad\|\phi-m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}}\|_{L^{2}(\mathbb{R})}\leq\delta$
and
$E(\phi)-E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})$ $\geq-\|\nabla E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})\|_{L^{2}(\mathbb{R})}\cdot\delta-\frac{1}{2}\delta^{2}.$
Compared to (3.16), we have control of $\|\nabla E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})\|_{L^{2}(\mathbb{R})}$ thanks to the condition $d\leq\min_{i\neq j}|\xi_{i}^{*}-\xi_{j}^{*}|$, which implies
$E(\phi)-E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})$ $\geq-e^{-(1-\eta)d}\cdot\delta-\frac{1}{2}\delta^{2},$
where we used Lemma 3.6. Then, under the condition $\frac{1}{2}e^{-d}\geq c\cdot\delta$ and Lemma 3.12,
$E(\phi)$ $=\big(E(\phi)-E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})\big)+E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})$
$\geq\big(E(\phi)-E(m_{\xi_{1}^{*},\dots,\xi_{Q}^{*}})\big)+|Q|E_{\textup{kink}}+c\cdot e^{-(1+\eta)d}$
$\geq|Q|E_{\textup{kink}}+c\cdot e^{-(1+\eta)d}-e^{-(1-\eta)d}\cdot\delta-\frac{1}{2}\delta^{2}$
$\geq|Q|E_{\textup{kink}}+\frac{c}{2}\cdot e^{-(1+\eta)d}.$
This completes the proof.
@@ 4.1. Approximate multi-soliton manifold
$\overline{L}_{\varepsilon}:=L_{\varepsilon}-\varepsilon^{-\frac{1}{2}+2\eta}-1=L_{\varepsilon}(1-\varepsilon^{\eta}-\varepsilon^{\frac{1}{2}-\eta})\sim L_{\varepsilon}.$ (4.2)
@@
$\|m_{\xi}-m_{\xi}^{\varepsilon}\|_{L^{p}(\mathbb{R})}\lesssim e^{-c\varepsilon^{-\frac{1}{2}+2\eta}}$ (4.3)
$\|\partial_{x}m_{\xi}-\partial_{x}m_{\xi}^{\varepsilon}\|_{L^{p}(\mathbb{R})}\lesssim e^{-c\varepsilon^{-\frac{1}{2}+2\eta}}.$ (4.4)
@@
$m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}=\sum_{j=1}^{k}m^{\varepsilon}_{\xi_{j}},$ (4.5)
@@
$\|m_{\xi_{1},\dots,\xi_{k}}-m_{\xi_{1},\dots,\xi_{k}}^{\varepsilon}\|_{L^{p}(\mathbb{R})}\lesssim ke^{-c\varepsilon^{-\frac{1}{2}+2\eta}}$ (4.6)
$\|\partial_{x}m_{\xi_{1},\dots,\xi_{k}}-\partial_{x}m_{\xi_{1},\dots,\xi_{k}}^{\varepsilon}\|_{L^{p}(\mathbb{R})}\lesssim ke^{-c\varepsilon^{-\frac{1}{2}+2\eta}},$ (4.7)
@@
$\mathcal{M}_{k}^{\varepsilon}:=\big\{m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}:-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\leq\xi_{k}\leq\overline{L}_{\varepsilon}\big\},$ (4.8)
@@
$\mathcal{M}^{\varepsilon,<d}_{k}:=\Big\{m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}:-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\xi_{k}\leq\overline{L}_{\varepsilon}\;\;\textup{and}\;\;\min_{i\neq j}|\xi_{i}-\xi_{j}|<d\Big\},$ (4.9)
@@
$\mathcal{M}^{\varepsilon,\geq d}_{k}:$ $=\{m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}:-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\xi_{k}\leq\overline{L}_{\varepsilon}\quad\textup{and}\quad\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d\},$ (4.10)
@@
4.2. Tangent and normal spaces of the manifold and disintegration formula
@@
Remark 4.1.
Notice that the multi-soliton manifold $\mathcal{M}_{k}^{\varepsilon}$ and the collision manifold $\mathcal{M}_{k}^{\varepsilon,<d}$, defined in (4.8) and (4.9), are topological manifolds but fail to be differentiable manifolds for $k\geq 3$ due to the presence of collision strata.
For example, when $k=3$, the parameter space $\{\xi_{1}\leq\xi_{2}\leq\xi_{3}\}$ has two codimension-one faces, given by $\xi_{1}=\xi_{2}$ and $\xi_{2}=\xi_{3}$. Each point on these faces admits a neighborhood that is locally diffeomorphic to $\mathbb{R}_{+}\times\mathbb{R}^{2}$. However, there is also an edge corresponding to $\xi_{1}=\xi_{2}=\xi_{3}$ and each point on this edge has a neighborhood that is locally diffeomorphic to $\mathbb{R}^{2}_{+}\times\mathbb{R}$. Since $\mathbb{R}^{2}_{+}$ is the quadrant with a corner, which is homeomorphic but not diffeomorphic to the half-plane, $\mathcal{M}_{k}$ and $\mathcal{M}_{k}^{\varepsilon,<d}$ fail to be a differentiable manifold (or a differentiable manifold with boundary), although they are topological manifolds with boundaries. For $k>3$, the geometry is more complicated, as $\mathcal{M}_{k}^{\varepsilon}$ and $\mathcal{M}_{k}^{\varepsilon,<d}$ decompose into strata of different codimensions. In particular, understanding the structure of the collision manifold is of independent interest from a geometric perspective.
Therefore, in order to carry out Riemannian geometric considerations such as defining tangent and normal vectors, we restrict our analysis to the smooth non-collision manifold $\mathcal{M}_{k}^{\varepsilon,\geq d}$ by removing the collision region $\min_{i\neq j}|\xi_{i}-\xi_{j}|<d$.
@@
$V_{\xi_{1},\dots,\xi_{k}}=\big\{v\in L^{2}:\;\langle v,\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}\rangle=0\quad\mbox{for all }1\leq j\leq k\big\}.$ (4.11)
@@
$|\langle\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}},\partial_{\xi_{i}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}\rangle|\lesssim e^{-c|\xi_{i}-\xi_{j}|}$ (4.12)
@@
We are now ready to define the projection map $\pi^{\varepsilon}$ onto the multi-soliton manifold, introduced in Theorem 1.4. Let $\mathcal{M}$ be a compact manifold in a Hilbert space $\mathcal{H}$. If $\delta>0$ is sufficiently small, we can assign to any $\phi\in\mathcal{H}$ with $\text{dist}(\phi,\mathcal{M})<\delta$ a unique closest point $\pi(\phi)$ in the manifold $\mathcal{M}$. This follows from the $\varepsilon$-neighborhood theorem [24, p.69]. If $\text{dist}(\phi,\mathcal{M})\geq\delta$, then we set $\pi(\phi)=0$. Recall that
$\mathcal{M}^{\varepsilon,\geq d}_{k}:$ $=\{m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}:-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\xi_{k}\leq\overline{L}_{\varepsilon}\quad\textup{and}\quad\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d\}$
is a $k$-dimensional compact manifold. Therefore, if a field $\phi$ satisfies
$\text{dist}(\phi,\mathcal{M}^{\varepsilon,\geq d}_{k})=\inf_{\begin{subarray}{c}-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\xi_{k}\leq\overline{L}_{\varepsilon}\\ \min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d\end{subarray}}\|\phi-m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}\|_{L^{2}}<\delta$ (4.13)
@@
$\pi^{\varepsilon}(\phi)=\begin{cases}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}},\;&\text{dist}(\phi,\mathcal{M}^{\varepsilon,\geq d}_{k})<\delta\\ 0,\;&\text{dist}(\phi,\mathcal{M}^{\varepsilon,\geq d}_{k})\geq\delta.\end{cases}$ (4.15)
@@
$\hphantom{XXXXXX}\cdot\textup{Det}_{\xi_{1},\dots,\xi_{k}}(v)\,\mu^{\perp}_{\varepsilon,\xi_{1},\dots,\xi_{k}}(dv)\,d\sigma(\xi_{1},\dots,\xi_{k}),$ (4.16)
@@
$\textup{Det}_{\xi_{1},\dots,\xi_{k}}(v):=\det\big(\operatorname{Id}-W_{\xi_{1},\dots\xi_{k},v}\big),$ (4.17)
@@
The disintegration formula has been widely employed in recent works [33, 36, 22] in settings where the energy functional admits explicit minimizers and the analysis is carried out around a single-soliton manifold. However, to the best of our knowledge, in Proposition 7.4 this is the first time the disintegration formula is applied in a regime where no minimizers exist and the analysis is performed around a multi-soliton manifold.
@@
$t_{j}:=\frac{w_{j}}{\|w\|_{L^{2}}}\qquad\textup{and}\qquad w_{j}:=\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}-\sum_{\ell<j}\langle\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}},t_{\ell}\rangle_{L^{2}}\;t_{\ell}.$ (4.18)
@@
$d\sigma(\xi_{1},\dots,\xi_{k})=|\gamma(\xi_{1},\dots,\xi_{k})|d\xi_{1}\dots d\xi_{k},$ (4.19)
@@
$|\gamma(\xi_{1},\dots,\xi_{k})|=k\|\partial_{x}m\|_{L^{2}(\mathbb{R})}^{2}(1+O(e^{-cd})),$
@@
$W_{\xi_{1},\dots,\xi_{k},v}:T_{\xi_{1},\dots,\xi_{k}}\mathcal{M}^{\varepsilon,\geq d}_{k}\to T_{\xi_{1},\dots,\xi_{k}}\mathcal{M}^{\varepsilon,\geq d}_{k},$ (4.20)
where $T_{\xi_{1},\dots,\xi_{k}}\mathcal{M}^{\varepsilon,\geq d}_{k}:=\text{span}\big\{t_{1},\dots,t_{k}\big\}$ is the tangent space of $\mathcal{M}^{\varepsilon,\geq d}_{k}$ at $m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}$ and $t_{1},\dots,t_{k}$ are obtained from the Gram–Schmidt orthonormalization described above (see (4.18)). Specifically, the Weingarten map $W_{\xi_{1},\dots,\xi_{k},v}$ in the basis $\{t_{1},\dots,t_{k}\}$ is given by
$\begin{pmatrix}\langle-\partial_{\xi_{1}}N_{\xi_{1},\dots,\xi_{k}}(v),t_{1}\rangle_{L^{2}}&\cdots&\langle-\partial_{\xi_{1}}N_{\xi_{1},\dots,\xi_{k}}(v),t_{k}\rangle_{L^{2}}\\ \vdots&\ddots&\vdots\\ \langle-\partial_{\xi_{k}}N_{\xi_{1},\dots\xi_{k}}(v),t_{1}\rangle_{L^{2}}&\cdots&\langle-\partial_{\xi_{k}}N_{\xi_{1},\dots\xi_{k}}(v),t_{k}\rangle_{L^{2}}\end{pmatrix}$,
where
$L^{2}\ni v\mapsto N_{\xi_{1},\dots,\xi_{k}}(v)$
is a parametrization of the normal space to $\mathcal{M}^{\varepsilon,\geq d}_{k}$ at a point $m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}$. In particular, the $k\times k$ determinant
$\det\big(\operatorname{Id}-W_{\xi_{1},\dots,\xi_{k},v}\big)=1+O\big(\|v\|_{L^{\infty}}^{k}\big)$ (4.21)
@@
as $\min_{j\neq\ell}|\xi_{j}-\xi_{\ell}|\geq d\to\infty$. ∎

### v2

@@ 3.3. Collision regime
$=-\frac{1}{2}\sum_{e}\Big(\cos((1-e_{1})m_{i_{1}}+\cdots+(1-e_{n})m_{i_{n}})$
$\qquad\qquad\qquad+\cos((1+e_{1})m_{i_{1}}+\cdots+(1+e_{n})m_{i_{n}})\Big)\prod_{j=1}^{n}e_{j}$
@@ 4.1. Approximate multi-soliton manifold
(4.2) $\overline{L}_{\varepsilon}:=L_{\varepsilon}-\varepsilon^{-\frac{1}{2}+2\eta}-1=L_{\varepsilon}(1-\varepsilon^{\eta}-\varepsilon^{\frac{1}{2}-\eta})\sim L_{\varepsilon}.$
@@
(4.3) $\|m_{\xi}-m_{\xi}^{\varepsilon}\|_{L^{p}(\mathbb{R})}\lesssim e^{-c\varepsilon^{-\frac{1}{2}+2\eta}}$
(4.4) $\|\partial_{x}m_{\xi}-\partial_{x}m_{\xi}^{\varepsilon}\|_{L^{p}(\mathbb{R})}\lesssim e^{-c\varepsilon^{-\frac{1}{2}+2\eta}}.$
@@
(4.5) $m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}=\sum_{j=1}^{k}m^{\varepsilon}_{\xi_{j}},$
@@
(4.6) $\|m_{\xi_{1},\dots,\xi_{k}}-m_{\xi_{1},\dots,\xi_{k}}^{\varepsilon}\|_{L^{p}(\mathbb{R})}\lesssim ke^{-c\varepsilon^{-\frac{1}{2}+2\eta}}$
(4.7) $\|\partial_{x}m_{\xi_{1},\dots,\xi_{k}}-\partial_{x}m_{\xi_{1},\dots,\xi_{k}}^{\varepsilon}\|_{L^{p}(\mathbb{R})}\lesssim ke^{-c\varepsilon^{-\frac{1}{2}+2\eta}},$
@@
(4.8) $\mathcal{M}_{k}^{\varepsilon}:=\big\{m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}:-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\leq\xi_{k}\leq\overline{L}_{\varepsilon}\big\},$
@@
(4.9) $\mathcal{M}^{\varepsilon,<d}_{k}:=\Big\{m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}:-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\xi_{k}\leq\overline{L}_{\varepsilon}\;\;\textup{and}\;\;\min_{i\neq j}|\xi_{i}-\xi_{j}|<d\Big\},$
@@
(4.10) $\mathcal{M}^{\varepsilon,\geq d}_{k}:$ $=\{m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}:-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\xi_{k}\leq\overline{L}_{\varepsilon}\quad\textup{and}\quad\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d\},$
@@
4.2. Tangent and normal spaces and disintegration formula
@@
When $k>1$, at the collisions, $\mathcal{M}^{\varepsilon}_{k}$ will have singularities and thus lose the smooth manifold structure. We postpone this discussion to the end of this section (see Remark 4.5 and the discussion above). Therefore, in order to carry out Riemannian geometric considerations such as defining tangent and normal vectors, we restrict our analysis to the smooth non-collision manifold $\mathcal{M}_{k}^{\varepsilon,\geq d}$ by removing the collision region.
@@
(4.11) $V_{\xi_{1},\dots,\xi_{k}}=\big\{v\in L^{2}:\;\langle v,\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}\rangle=0\quad\mbox{for all }1\leq j\leq k\big\}.$
@@
(4.12) $|\langle\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}},\partial_{\xi_{i}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}\rangle|\lesssim e^{-c|\xi_{i}-\xi_{j}|}$
@@
We are now ready to define the projection map $\pi^{\varepsilon}$ onto the multi-soliton manifold, introduced in Theorem 1.4. Let $\mathcal{M}$ be a compact manifold in a Hilbert space $\mathcal{H}$. If $\delta>0$ is sufficiently small, we can assign to any $\phi\in\mathcal{H}$ with $\text{dist}(\phi,\mathcal{M})<\delta$ a unique closest point $\pi(\phi)$ in the manifold $\mathcal{M}$. This follows from the $\varepsilon$-neighborhood theorem [24, p.69]. If $\text{dist}(\phi,\mathcal{M})\geq\delta$, then we set $\pi(\phi)=0$. Recall that $\mathcal{M}^{\varepsilon,\geq d}_{k}$ defined in (4.10) is a $k$-dimensional compact manifold. Therefore, if a field $\phi$ satisfies
(4.13) $\text{dist}(\phi,\mathcal{M}^{\varepsilon,\geq d}_{k})=\inf_{\begin{subarray}{c}-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\xi_{k}\leq\overline{L}_{\varepsilon}\\ \min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d\end{subarray}}\|\phi-m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}\|_{L^{2}}<\delta$
@@
(4.15) $\pi^{\varepsilon}(\phi)=\begin{cases}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}},\;&\text{dist}(\phi,\mathcal{M}^{\varepsilon,\geq d}_{k})<\delta\\ 0,\;&\text{dist}(\phi,\mathcal{M}^{\varepsilon,\geq d}_{k})\geq\delta.\end{cases}$
@@
(4.16) $\hphantom{XXXXXX}\cdot\textup{Det}_{\xi_{1},\dots,\xi_{k}}(v)\,\mu^{\perp}_{\varepsilon,\xi_{1},\dots,\xi_{k}}(dv)\,d\sigma(\xi_{1},\dots,\xi_{k}),$
@@
(4.17) $\textup{Det}_{\xi_{1},\dots,\xi_{k}}(v):=\det\big(\operatorname{Id}-W_{\xi_{1},\dots\xi_{k},v}\big),$
@@
The disintegration formula has been widely employed in recent works [33, 36, 22] in settings where the energy functional admits explicit minimizers and the analysis is carried out around a single-soliton manifold. However, to the best of our knowledge, this is the first time the disintegration formula is applied in a regime where no minimizers exist and the analysis is performed around a multi-soliton manifold. See also Proposition 7.2.
@@
(4.18) $t_{j}:=\frac{w_{j}}{\|w\|_{L^{2}}}\qquad\textup{and}\qquad w_{j}:=\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}-\sum_{\ell<j}\langle\partial_{\xi_{j}}m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}},t_{\ell}\rangle_{L^{2}}\;t_{\ell}.$
@@
(4.19) $d\sigma(\xi_{1},\dots,\xi_{k})=|\gamma(\xi_{1},\dots,\xi_{k})|d\xi_{1}\dots d\xi_{k},$
@@
(4.20) $|\gamma(\xi_{1},\dots,\xi_{k})|=k\|\partial_{x}m\|_{L^{2}(\mathbb{R})}^{2}(1+O(e^{-cd})),$
@@
(4.21) $W_{\xi_{1},\dots,\xi_{k},v}:T_{\xi_{1},\dots,\xi_{k}}\mathcal{M}^{\varepsilon,\geq d}_{k}\to T_{\xi_{1},\dots,\xi_{k}}\mathcal{M}^{\varepsilon,\geq d}_{k},$
where $T_{\xi_{1},\dots,\xi_{k}}\mathcal{M}^{\varepsilon,\geq d}_{k}:=\text{span}\big\{t_{1},\dots,t_{k}\big\}$ is the tangent space of $\mathcal{M}^{\varepsilon,\geq d}_{k}$ at $m^{\varepsilon}_{\xi_{1},\dots,\xi_{k}}$. Specifically, the Weingarten map $W_{\xi_{1},\dots,\xi_{k},v}$ in the basis $\{t_{1},\dots,t_{k}\}$ is given by
$\begin{pmatrix}\langle-\partial_{\xi_{1}}N_{\xi_{1},\dots,\xi_{k}}(v),t_{1}\rangle_{L^{2}}&\cdots&\langle-\partial_{\xi_{1}}N_{\xi_{1},\dots,\xi_{k}}(v),t_{k}\rangle_{L^{2}}\\ \vdots&\ddots&\vdots\\ \langle-\partial_{\xi_{k}}N_{\xi_{1},\dots\xi_{k}}(v),t_{1}\rangle_{L^{2}}&\cdots&\langle-\partial_{\xi_{k}}N_{\xi_{1},\dots\xi_{k}}(v),t_{k}\rangle_{L^{2}}\end{pmatrix}$.
In particular, the $k\times k$ determinant
(4.22) $\det\big(\operatorname{Id}-W_{\xi_{1},\dots,\xi_{k},v}\big)=1+O\big(\|v\|_{L^{\infty}}^{k}\big)$
@@
as $\min_{j\neq\ell}|\xi_{j}-\xi_{\ell}|\geq d\to\infty$.
We conclude this section with some discussion on the singular geometry in the collision regime, since understanding the structure of the collision manifold is of independent interest from a geometric perspective.
When $k=2$, as $|\xi_{1}-\xi_{2}|$ becomes small, we lose the almost orthogonality (4.12) of the tangent vectors $\partial_{\xi_{1}}m^{\varepsilon}_{\xi_{1},\xi_{2}}$ and $\partial_{\xi_{2}}m^{\varepsilon}_{\xi_{1},\xi_{2}}$. The Jacobian decreases rank by $1$ when $\xi_{1}=\xi_{2}$. Writing in the coordinates $\xi_{1}=\xi-s$ and $\xi_{2}=\xi+s$ where $\xi$ is the “center of mass”, the differentiation of the map $(\xi_{1},\xi_{2})\mapsto m^{\varepsilon}_{\xi_{1},\xi_{2}}\in L^{2}$ in the $s$-direction vanishes at the boundary.
Refer to caption    $\rightarrow$    Refer to caption
Figure 3. A map $\{(\xi_{1},\xi_{2}):\xi_{1}<\xi_{2}\}\to\mathbb{R}^{2}$ where tangent vectors are almost orthogonal in the bulk, but asymptotically collinear near the boundary. The two families of curves in the bulk are lines of constant $\xi_{1}$ and $\xi_{2}$. Picture drawn with a toy model $(\xi_{1},\xi_{2})\mapsto(\xi_{1}+\xi_{2},\xi_{1}^{2}+\xi_{2}^{2})$ to illustrate $(\xi_{1},\xi_{2})\mapsto m_{\xi_{1},\xi_{2}}$ which is not drawable in infinite dimensions.
Remark 4.5.
Although for $k=2$ one may endow another differentiable structure to turn $\mathcal{M}_{2}^{\varepsilon}$ and $\mathcal{M}_{2}^{\varepsilon,<d}$ into smooth manifolds with boundaries, in general $\mathcal{M}_{k}^{\varepsilon}$ and $\mathcal{M}_{k}^{\varepsilon,<d}$ are just topological manifolds and fail to be differentiable manifolds for $k\geq 3$ due to the presence of collision strata.
For example, when $k=3$, the parameter space $\{\xi_{1}\leq\xi_{2}\leq\xi_{3}\}$ has two codimension-one faces, given by $\xi_{1}=\xi_{2}$ and $\xi_{2}=\xi_{3}$. Each point on these faces admits a neighborhood that is locally diffeomorphic to $\mathbb{R}_{+}\times\mathbb{R}^{2}$. However, there is also an edge corresponding to $\xi_{1}=\xi_{2}=\xi_{3}$ and each point on this edge has a neighborhood that is locally diffeomorphic to $\mathbb{R}^{2}_{+}\times\mathbb{R}$. Since $\mathbb{R}^{2}_{+}$ is the quadrant with a corner, which is homeomorphic but not diffeomorphic to the half-plane, $\mathcal{M}_{k}$ and $\mathcal{M}_{k}^{\varepsilon,<d}$ fail to be a differentiable manifold (or a differentiable manifold with boundary), although they are topological manifolds with boundaries. For $k>3$, the geometry is more complicated, as $\mathcal{M}_{k}^{\varepsilon}$ and $\mathcal{M}_{k}^{\varepsilon,<d}$ decompose into strata of different codimensions.